OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)?

OpenStudy (anonymous):

Slope = \(\frac{y_2-y_1}{x_2-x_2}\) = (14-4)/(4-2) = 10 / 2 = ... ?

OpenStudy (anonymous):

5 or 1/5?